OpenStudy (anonymous):

cubic function? help. I'm confused on the shrinking/stretching ):

OpenStudy (anonymous):

im gonna post an example D:

OpenStudy (anonymous):

\(y=ax^3+bx^2+cx+d\) if \(|a|>1\) it is a vertical stretch. if \(0<|a|<1\) then it is a vertical shrink

OpenStudy (anonymous):

is your function \(y=2(x-4)^3-1\)?

OpenStudy (anonymous):

im not sure about the the 2, bust (x-4)^3-1 i agree with

OpenStudy (anonymous):

maybe \(y=\frac{1}{2}(x-4)^3-1 \) you can determine \(a\) by using a point on the graph.

OpenStudy (anonymous):

\[1=a(6-4)^2-1 \Rightarrow 2=4a \Rightarrow a = \frac{1}{2}\]

OpenStudy (anonymous):

so it is a ________?

OpenStudy (anonymous):

how come it's a(6-4)^2? why is it squared. shouldn't it be cubed?

OpenStudy (anonymous):

@pgpilot326

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

cause you are substituting for a quadratic instead of a cube

OpenStudy (aum):

Can you type the parent function and the transformed function? It is not clear from the pic which is posted sideways.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

parent function is y=x^3 transformed= (answer) is 1/2(x-4)^3-2

OpenStudy (aum):

!. x in the parent function is replaced by (x-4). This translates the parent function to the RIGHT by 4 units. 2. Then (x-4)^3 is multiplied by 1/2. This reduces or shrinks the y-value by a factor of two. So there is vertical compression to half the original value. 3. Then subtraction of 2 causes a vertical translation DOWN by 2 units.

OpenStudy (anonymous):

@aum im doing the reverse way. How would i find the compression/stretch if given the point of (6,-1)?

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

how do i do the substitution, cause the guy above substitute the point in a quadratic function. But his answer was correct

OpenStudy (aum):

Are you given a graph from which you have to find the function? The link is sideways making it hard to read what is given and what is asked.

OpenStudy (anonymous):

type the transformed function into https://www.desmos.com/calculator

OpenStudy (anonymous):

yes im given a graph

OpenStudy (aum):

Can you post the pic upright where it is clear what is given and what is asked?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me rotate it

OpenStudy (anonymous):

oh nevermind

OpenStudy (aum):

Still I don't see anything that say the parent function is y = x^3.

OpenStudy (anonymous):

you have to identify it

OpenStudy (anonymous):

it was originally a cubic function, now it's transformed

OpenStudy (aum):

Okay, it does look like a cubic function shifted to the right. The flat part of the curve in the graph is at (4,-1) y = x^3 has the flat part at (0,0) So this function has been shifted to the right by 4 unit and down by 1 unit. Therefore, y = (x - 4)^3 - 1. And ....

OpenStudy (aum):

We see a point (6, 1) on the graph. That is, when x = 6, y = 1. Let us see what the equation y = (x - 4)^3 - 1 gives us when x = 6: y = (6 - 4)^3 - 1 = 8 - 1 = 7. So the y value instead of being 7 is 1. So there is a vertical compression. So y = a(x-4)^3 - 1. Where 'a' is the vertical compression/stretch. Now put x = 6 and y = 1 and solve for a. 1 = a(6 - 4)^3 - 1 1 = 8a - 1 8a = 2 a = 1/4 y = 1/4(x - 4)^3 - 1

OpenStudy (aum):

\[ \text{Parent function } ~ y = x^3~ \text{ transformed to: }~~y = \frac 14(x-4)^3 - 1 \]

OpenStudy (anonymous):

ah, oh the guy above before confused me

OpenStudy (anonymous):

ty i understand :D

OpenStudy (aum):

You are welcome.